# Drive Anytime Anywhere in Any Condition

### v1

@@ 1 Introduction
The evolution of vision-centric autonomous driving hinges critically on robust perception systems that interpret and respond accurately to complex and dynamic environments [3, 15, 72]. As shown in Fig. 1, central to these systems are four mainstream perception tasks: bird’s eye view (BEV) detection, HD map segmentation, semantic occupancy prediction, and multi-view depth estimation [57]. Mastery of these tasks ensures that vehicles can navigate safely and interact intelligently with their surroundings [45, 25, 82].
@@
Autonomous vehicles must operate reliably under a variety of conditions, many of which may deviate significantly from the environments seen during the training of their perception systems [80]. These out-of-distribution (OoD) scenarios, such as adverse weather, sensor malfunctions, and unpredictable environmental changes, pose significant challenges [19]. They can severely impact the performance of perception systems, leading to degraded situational awareness and potentially unsafe driving decisions [29, 35, 2].
@@ 2.3 Semantic Occupancy Prediction
Semantic occupancy prediction aims to predict the occupancy and semantic labels of 3D scenes, unlike occupancy grid mapping [61, 73], which generates probabilistic maps from range measurements and assumes static scenes. This task does not require range sensors, making it more versatile. Semantic Scene Completion (SSC) closely relates to semantic occupancy prediction, inferring dense 3D occupancy with semantic labels. MonoScene [4] employs a U-Net for this from monocular RGB images. TPVFormer [26] extends to multi-camera input but lacks dense supervision. Recently, improved pipelines are now able to generate dense occupancy ground truth for better training [79, 74]. Methods like VoxFormer [47] and OccDepth [59] use depth-aware techniques to further enhance the performance. Another task that is closely related to SSC is LiDAR segmentation [60, 32, 52, 54]. Sharing a similar objective, the LiDAR segmentation models aim to segment the holistic 3D scenes acquired by the LiDAR sensor [7, 33, 28, 36]. It is worth mentioning that the SSC labels are often obtained by concatenating multi-frame semantic LiDAR point clouds [5, 22, 46].
@@ 2.4 Multi-View Depth Estimation
2.5 Out of Distribution Robustness
The resilience of models against unexpected disturbances, i.e., out-of-distribution (OoD), is a critical area of focus in machine learning, particularly under conditions that deviate from typical training environments. Several benchmarks such as ImageNet-C [19], ObjectNet [1], and ImageNetV2 [66] have been established to systematically evaluate the robustness of 2D image classifiers against a diverse range of corruptions. For example, ImageNet-C [19] introduces controlled perturbations to pristine ImageNet [8] samples, including compression artifacts and motion blur, to test classifier resilience. Conversely, ObjectNet [1] challenges models with a test set characterized by significant variations in object rotation, background, and viewpoint, emphasizing the importance of adaptability in real-world applications. Hendrycks et al. [21] highlight the direct correlation between a model’s ability to withstand synthetic corruptions and its performance in practical settings. In the realm of 3D perception, recent initiatives have aimed to extend these concepts to more complex models. RoboDepth [35, 31] establishes a robustness benchmark for monocular depth estimation under various corruptions. Robo3D [29] aims to probe the robustness of LiDAR-based perception models in adverse weather conditions and sensor malfunctions. RoboBEV [80] and MetaBEV propose to understand the BEV detection model’s robustness under external distributions. Most recently, MultiCorrupt [2] combines several 2D and 3D corruptions together to further analyze the robustness of multi-modal perception models.
@@ 3.3 Challenge Phases
  • Phase 1: preliminary development, from January to March.
@@ 7 Appendix
In this appendix, we summarize the information of the challenge organizers, technical committee members, and challenge participants from each track.

### v2

@@ 1 Introduction
The evolution of vision-centric autonomous driving hinges critically on robust perception systems that interpret and respond accurately to complex and dynamic environments [3, 15, 74]. As shown in Fig. 1, central to these systems are four mainstream perception tasks: bird’s eye view (BEV) detection, HD map segmentation, semantic occupancy prediction, and multi-view depth estimation [59]. Mastery of these tasks ensures that vehicles can navigate safely and interact intelligently with their surroundings [44, 46, 25, 85].
@@
Autonomous vehicles must operate reliably under a variety of conditions, many of which may deviate significantly from the environments seen during the training of their perception systems [82, 83]. These out-of-distribution (OoD) scenarios, such as adverse weather, sensor malfunctions, and unpredictable environmental changes, pose significant challenges [19]. They can severely impact the performance of perception systems, leading to degraded situational awareness and potentially unsafe driving decisions [29, 35, 2].
@@ 2.3 Semantic Occupancy Prediction
Semantic occupancy prediction aims to predict the occupancy and semantic labels of 3D scenes, unlike occupancy grid mapping [63, 75], which generates probabilistic maps from range measurements and assumes static scenes. This task does not require range sensors, making it more versatile. Semantic Scene Completion (SSC) closely relates to semantic occupancy prediction, inferring dense 3D occupancy with semantic labels. MonoScene [4] employs a U-Net for this from monocular RGB images. TPVFormer [26] extends to multi-camera input but lacks dense supervision. Recently, improved pipelines are now able to generate dense occupancy ground truth for better training [81, 76]. Methods like VoxFormer [48] and OccDepth [61] use depth-aware techniques to further enhance the performance. Another task that is closely related to SSC is LiDAR semantic segmentation [62, 32, 53, 55, 56]. Sharing a similar objective, the LiDAR segmentation models aim to segment the holistic 3D scenes acquired by the LiDAR sensor [7, 33, 28, 36]. It is worth mentioning that the SSC labels are often obtained by concatenating multi-frame semantic LiDAR point clouds [5, 22, 47].
@@ 2.4 Multi-View Depth Estimation
2.5 Out-of-Distribution Robustness
The resilience of models against unexpected disturbances, i.e., out-of-distribution (OoD), is a critical area of focus in machine learning, particularly under conditions that deviate from typical training environments. Several benchmarks such as ImageNet-C [19], ObjectNet [1], and ImageNetV2 [68] have been established to systematically evaluate the robustness of 2D image classifiers against a diverse range of corruptions. For example, ImageNet-C [19] introduces controlled perturbations to pristine ImageNet [8] samples, including compression artifacts and motion blur, to test classifier resilience. Conversely, ObjectNet [1] challenges models with a test set characterized by significant variations in object rotation, background, and viewpoint, emphasizing the importance of adaptability in real-world applications. Hendrycks et al. [21] highlight the direct correlation between a model’s ability to withstand synthetic corruptions and its performance in practical settings. In the realm of 3D perception, recent initiatives have aimed to extend these concepts to more complex models. RoboDepth [35, 31] establishes a robustness benchmark for monocular depth estimation under various corruptions. Robo3D [29] aims to probe the robustness of LiDAR-based perception models in adverse weather conditions and sensor malfunctions. RoboBEV [82, 83] and MetaBEV propose to understand the BEV detection model’s robustness under external distributions. Most recently, MultiCorrupt [2] combines several 2D and 3D corruptions together to further analyze the robustness of multi-modal perception models.
@@ 3.3 Challenge Phases
  • Phase 1: Preliminary exploration, from January to March.
@@ 7 Appendix
In this appendix, we summarize the public resources of this challenge (cf. Sec. 7.1), and the information of 1) the challenge organizers (cf. Sec. 7.2), 2) the technical committee members (cf. Sec. 7.3), as well as 3) the challenge participants from each track (cf. Sec. 7.4). We also include the track report from each winning team (cf. Sec. 7.5).
7.1 Public Resources
@@ 7.4.5 Track 5: Robust Multi-Modal BEV Detection
7.5 Track Reports
  • Xu Cao, Hao Lu, and Ying-Cong Chen. “Towards Robust Multi-Camera 3D Object Detection through Temporal Sequence Mix Augmentation”, Technical Report, 2024.
  • Caixin Kang, Xinning Zhou, Chengyang Ying, Wentao Shang, Xingxing Wei, and Yinpeng Dong. “MVE: Multi-View Enhancer for Robust Bird’s Eye View Object Detection”, Technical Report, 2024.
  • Bo Yang, Shengyin Jiang, Zeliang Ma, Dengyi Ji, and Haiwen Li. “FocalAngle3D: An Angle-Enhanced Two-Stage Model for 3D Detection”, Technical Report, 2024.
  • Xingliang Huang and Yu Tian. “Models and Data Enhancements for Robust Map Segmentation in Autonomous Driving”, Technical Report, 2024.
  • Xiaoshuai Hao, Yifan Yang, Hui Zhang, Mengchuan Wei, Yi Zhou, Haimei Zhao, and Jing Zhang. “Using Temporal Information and Mixing-Based Data Augmentations for Robust HD Map Construction”, Technical Report, 2024.
  • Genghua Kou, Fan Jia, Yingfei Liu, Tiancai Wang, and Ying Li. “MultiViewRobust: Scaling Up Pretrained Models for Robust Map Segmentation”, Technical Report, 2024.
  • Jinke Li, Xiao He, and Xiaoqiang Cheng. “ViewFormer: Spatiotemporal Modeling for Robust Occupancy Prediction”, Technical Report, 2024.
  • Bingyang Zhang, Lirong Zhao, Dianlei Ding, Fangsheng Liu, Yixiang Yan, and Hongming Wang. “Robust Occupancy Prediction based on Enhanced SurroundOcc”, Technical Report, 2024.
  • Nanfei Ye, Lun Luo, Xun Wu, Yubo Tian, Zhe Cao, Yunfan Li, Yiwei Zuo, Wenjie Liu, and Yi Ren. “Improving Out-of-Distribution Robustness of Occupancy Prediction Networks with Advanced Loss Functions”, Technical Report, 2024.
  • Yifan Mao, Ming Li, Jian Liu, Jiayang Liu, Zihan Qin, Chunxi Chu, Jialei Xu, Wenbo Zhao, Junjun Jiang, and Xianming Liu. “DINO-SD for Robust Multi-View Supervised Depth Estimation”, Technical Report, 2024.
  • Ziyan Wang, Chiwei Li, Shilong Li, Chendong Yuan, Songyue Yang, Wentao Liu, Peng Chen, and Bin Zhou. “Fusing Features Across Scales: A Semi-Supervised Attention-Based Approach for Robust Depth Estimation”, Technical Report, 2024.
  • Yubo Wang, Chi Zhang, and Jianhang Sun. “SD-ViT: Performance and Robustness Enhancements of MonoViT for Multi-View Depth Estimation”, Technical Report, 2024.
  • Hai Chen, Xiao Yang, and Lizhong Wang. “ASF: Robust 3D Object Detection by Solving Sensor Failures”, Technical Report, 2024.
  • Caixin Kang, Xinning Zhou, Chengyang Ying, Wentao Shang, Xingxing Wei, and Yinpeng Dong. “Cross-Modal Transformers for Robust Multi-Modal BEV Detection”, Technical Report, 2024.
  • Dongyi Fu, Yongchun Lin, Huitong Yang, Haoang Li, Yadan Luo, Xianjing Cheng, and Yong Xu. “RobuAlign: Robust Alignment in Multi-Modal 3D Object Detection”, Technical Report, 2024.